Find the perimeter and area of a right triangle if one leg measures 12 yd. and the other leg measures 35 yd.
Answer by rwm(914)

12^2+35^2=c^2
c=37
perimeter=12+35+37=84
area=1/2 bh
area=1/2*12*35=6*35=210